OpenStudy (anonymous):

Carmela has a deck of 26 alphabet cards. She mixes the cards, places them face down on the table, and turns them over one at a time. What is the probability that the first two cards she turns over are vowel cards (a, e, i, o, u)?

OpenStudy (anonymous):

\[\frac{5}{26}\times \frac{4}{25}\]

OpenStudy (anonymous):

Do I do cross product?